What number increased by 20 percent equals 78 ?
A number that is 42% less than c.
150% of 50% of z
Thx,
kevin

Answer by jorel1380(3719)

1.2n=78
n=78/1.2=65
...........
c-.42c=c(1-.42)=.58c
............
1.5(.5)(z)=0.75z
☺☺☺☺